Assume that today is December 31, 2019, and that the following information applies to Abner Airlines: After-tax operating income [EBIT(1 - T)] for 2020 is expected to be $500 million.

The depreciation expense for 2020 is expected to be $170 million.

The capital expenditures for 2020 are expected to be $225 million.

No change is expected in net operating working capital.

The free cash flow is expected to grow at a constant rate of 3% per year.

The required return on equity is 14%.

The WACC is 12%.

The firm has $191 million of non-operating assets.

The market value of the company's debt is $2.686 billion.

290 million shares of stock are outstanding.

Using the corporate valuation model approach, what should be the company's stock price today? Do not round intermediate calculations. Round your answer to the nearest cent. $
Answer #1

FCFF = EBIT x (1 - T) + Depreciation - NWC - Capex

= 500 + 170 - 225

= $445 million

Value of firm = FCFF / (WACC - g) = 445 / (12% - 3%) = $4,944.44 million

Value of firm = Equity + Debt - Non-op assets

=> Equity = 4,944.44 - 2,686 + 191 = $2,449.44 million

Stock Price = Equity Value / No. of shares = 2,449.44 / 290 = $8.45